# Optimal Methods for Unknown Piecewise Smooth Problems I: Convex Optimization

### v1

@@ 1 Introduction
$f^{*}:=\min_{x\in X}f(x).$ (1)
Under the assumption that the objective $f$ satisfies the quadratic growth (QG) condition (see Definition 2.2), we investigate the oracle complexity of first-order methods for finding an $\epsilon$-optimal solution, defined as the theoretical number of (sub)-gradient evaluations required to reach a point $x$ such that $f(x)-f^{*}\leq\epsilon$. We focus on unknown piecewise smooth (PWS) functions: the domain $X$ admits a finite partition into regions on each of which the function is smooth, but the partition is not known a priori (Definition 2.1). Such functions arise, for example, in quadratic Model Predictive Control (MPC) [42] and two-stage stochastic linear programming (SLP) [16], where those regions are determined by the active constraints and the optimal dual solution in the second-stage problem, respectively, and hence are not available a priori. They also arise in ReLU networks [59], where regions are determined by the composition of activation functions. As network depth increases, these pieces become more difficult to characterize a priori.
@@
$\hat{L}_{k}=\frac{2\left[f(x^{k+1})-f(x^{k})-\left\langle f^{\prime}(x^{k}),x^{k+1}-x^{k}\right\rangle\right]}{\left\|x^{k+1}-x^{k}\right\|^{2}}.$ (2)
@@
Example 1.1: Average vs. worst-case empirical Lipschitz smoothness constant In the PWS setting, the gradient is continuous within each smooth region but may be discontinuous across their boundaries. This discontinuity can cause the empirical Lipschitz smoothness constant $\hat{L}_{k}$ (2) to spike when iterates $x^{k}$ and $x^{k+1}$ belong to different regions. Since the worst-case empirical Lipschitz smoothness constant $L_{\max}$ is determined by these spikes, it can be significantly larger than the average $L_{\text{avg}}$. To illustrate this, consider the iterates generated by applying the Polyak step size to the following one-dimensional objective function with two smooth regions, $\{x\geq 0\}$ and $\{x<0\}$: $\min_{x\in\left[-M,+\infty\right]}f(x;i)=(ix+1)\cdot\mathbf{1}_{\left\{x\geq 0\right\}}+e^{x}\cdot\mathbf{1}_{\left\{x<0\right\}}.$ The objective is parameterized by $i\in\mathbb{N}_{+}$ and $M\in\mathbb{R}$. Starting from an initial point $x^{0}=1/i$, we obtain $L_{\max}=i^{2}e^{-1/i}/2$, while $L_{\text{avg}}\rightarrow e^{-M}$ as the method progresses. Consequently, the ratio $L_{\max}/L_{\text{avg}}$ can be arbitrarily large for certain problem instances. Figure 2 illustrates how these two Lipschitz constants evolve over iterations.
@@
    Optimal convergence rate for PWS optimization. We introduce an accelerated bundle-level method termed Apex (Accelerated Prox-Level method for Exploring piecewise-smooth structure) in Section 3. We prove that with an adequate bundle of cutting planes, Apex achieves an optimal iteration complexity of $O\left(\kappa\sqrt{L/\mu^{*}}\log(1/\epsilon)\right)$ to compute an $\epsilon$-optimal solution for objectives satisfying $(\kappa,L)$-PWS (Definition 2.1) and $\mu^{*}$-QG (Definition 2.2) conditions. This result significantly improves upon the $O\left(\kappa^{2}L/\mu^{*}\log(1/\epsilon)\right)$ complexity reported in [74]. Our theoretical advancements translate to superior practical performance, as demonstrated in Figure 2. Furthermore, in Section 7, we establish the first lower complexity bound for PWS optimization, defining the theoretical benchmark for optimality. Consequently, Apex is the first algorithm to achieve an accelerated, provably optimal rate for piecewise-smooth optimization.
@@
    In Section 4, we introduce an accelerated bundle-level (BL) framework capable of constructing the W-certificate (see Definition 4.1) with significantly higher accuracy than previous methods [74]. While prior certification techniques could only bound the optimality gap within a factor of $O(L/\mu)$, our approach improves this bound to a constant factor, $O(1)$. This enhancement allows the certificate to serve as a robust proxy for the true optimality gap. Consequently, the certificate becomes a critical tool for dynamically updating the target level (an estimate of the optimal objective value, $f^{*}$), a essential parameter for the bundle-level method. Leveraging this accurate dynamic update, we design a new parameter-free algorithm, as detailed in our next contribution.
@@ 1.1 Related Work
$\frac{1}{2L}\left\|\nabla f\left(\bar{y}\right)\right\|^{2}\leq f\left(\bar{y}\right)-f^{*}\leq\frac{1}{2\mu}\left\|\nabla f\left(\bar{y}\right)\right\|^{2}\ .$ (3)
@@
For nonsmooth but Lipschitz functions, verifiability is preserved, but the strong guarantees in (3) no longer hold. By Rademacher’s theorem, Lipschitz functions are differentiable almost everywhere, which motivates defining the subdifferential as $\partial f(x):=\operatornamewithlimits{conv}\left\{g:g=\lim_{t\to\infty}f^{\prime}(x^{t}),x^{t}\to x\right\}$. With this definition, a point $x$ is stationary if $0\in\partial f(x)$. More generally, $x$ is a $(\delta,\varepsilon)$ near-approximate stationary point if $\left\|x-x^{\prime}\right\|\leq\varepsilon$ for some $\varepsilon$-stationary point $x^{\prime}$. While Tian and So [69] introduce the first oracle-polynomial-time algorithm for detecting near-approximate stationary points in piecewise affine functions, for general Lipschitz and bounded-below functions, finding such points cannot be guaranteed in polynomial time [41]. As a relaxation, recent work [72, 41] studies $(\delta,\varepsilon)$-Goldstein stationarity, which considers the convex hull of subgradients in a $\delta$-neighborhood. Given a finite sample $\mathcal{P}=\left\{x^{i}\right\}_{i=1}^{m}\subseteq\mathcal{B}(x;\delta)$, one can test $(\delta,\varepsilon)$-Goldstein stationarity by: $\min_{\lambda\in\Delta_{+}}\left\|\sum_{i=1}^{m}\lambda_{i}f^{\prime}(x^{i})\right\|$, where $\Delta_{+}$ is the probability simplex. Approximate Goldstein stationarity is thus computationally tractable, though it does not imply bounds on the optimality gap. On the other hand, for structured problems such as weakly convex functions, the Moreau stationary certificate is widely used. The Moreau envelope [58] is defined as: $f_{\lambda}(x):=\min_{y}\left\{f(y)+\lambda\left\|y-x\right\|^{2}/2\right\}$, which is smooth when $\lambda$ exceeds the weak convexity modulus, with $\nabla f_{\lambda}(x)=\lambda\left(x-x_{\lambda}\right)$ and $x_{\lambda}=\operatornamewithlimits{argmin}_{y}\left\{f(y)+\lambda\left\|y-x\right\|^{2}/2\right\}$. The main challenge is that computing $x_{\lambda}$ exactly is as hard as solving the original problem [19]. Approximating Moreau stationary [6] is possible, but they require an approximation $\hat{x}_{\lambda}$ with $\left\|\hat{x}_{\lambda}-{x}_{\lambda}\right\|$ sufficiently small, which in turn demands an additional termination certificate for the proximal subproblem. Additionally, Nemirovski et al. [60] developed an accuracy certificate framework for convex optimization, variational inequalities with monotone operators, and convex Nash equilibrium problems, showing how such certificates can be constructed via ellipsoid or cutting-plane methods. Nevertheless, this framework has two crucial limitations: it only applies to bounded domains, and it does not accommodate error bounds such as the QG property. As a result, it produces only sublinear convergence rates for certificate generation.
@@ 2 Preliminaries
Definition 2.1 ($(k,L)$-Piecewise smoothness, $\left(k,L\right)$-PWS).
@@
Definition 2.2 ($\mu^{*}$-Quadratic growth, $\mu^{*}$-QG).
@@
$\begin{aligned} f(x^{t+1})\leq&\ell_{f}(x^{t+1};x^{t})+{\hat{L}_{t}}\left\|x^{t+1}-x^{t}\right\|^{2}/2\\ \leq&f^{*}+{\hat{L}_{t}}(\left\|x^{t}-x^{*}\right\|^{2}-\left\|x^{t+1}-x^{*}\right\|^{2})/2\end{aligned}\ ,$ (4)
@@
$\begin{aligned} f(x^{t+i})\leq&\ell_{f}(x^{t+i};x^{t})+{\hat{L}_{t,i}}\left\|x^{t+i}-x^{t}\right\|^{2}/2\\ \leq&f^{*}+i\cdot{\hat{L}_{t,i}}\left(\left\|x^{t}-x^{*}\right\|^{2}-\left\|x^{t+i}-x^{*}\right\|^{2}\right)/2\end{aligned}\ ,\forall i\in\left[m\right]\ ,$ (5)
@@
$\begin{split}\underline{x}^{t}&=\left(1-\alpha_{t}\right)\bar{x}^{t-1}+\alpha_{t}x^{t-1}\\ x^{t}&=\operatornamewithlimits{argmin}_{x\in X}\left\|x-x^{t-1}\right\|^{2}\ ,\ \operatornamewithlimits{s.t.}\ \ell_{f}\left(x;\underline{x}^{i}\right)\leq\tilde{l}\ ,\ i\in\left[t\right]\ ,\\ \bar{x}^{t}&=\left(1-\alpha_{t}\right)\bar{x}^{t-1}+\alpha_{t}x^{t}\end{split}$
@@
$f(\bar{x}^{t})\leq$ $\ell_{f}(\bar{x}^{t};\underline{x}^{t})+{\hat{L}_{t}}\left\|\bar{x}^{t}-\underline{x}^{t}\right\|^{2}/2$ (6)
$\leq$ $f^{*}+{\hat{L}_{t}}\alpha_{t}^{2}\left(\left\|x^{t-1}-x^{*}\right\|^{2}-\left\|x^{t}-x^{*}\right\|^{2}\right)/2$
where $\hat{L}_{t}$ is the empirical Lipschitz smoothness constant determined between $\bar{x}^{t}$ and $\underline{x}^{t}$. A key structural relationship, $\bar{x}^{t}-\underline{x}^{t}=\alpha_{t}(x^{t}-x^{t-1})$, forms the backbone of the telescoping arguments in the convergence proofs of ABL and APL. However, this identity is valid only for adjacent pairs of iterates and fails to generalize to non-adjacent sequences, such as those arising in the BL method discussed in (5). This limitation becomes especially pronounced when analyzing PWS functions, where telescoping across non-adjacent points is required.
@@ 3 Accelerated Prox-level method for piecewise smooth optimization
1:Initialize: $\text{{Flag}}=\text{{\text{False}}}$
@@
$x^{t,i}\leftarrow\operatornamewithlimits{argmin}_{x\in X(t,i)}\left\|x-\bar{y}\right\|^{2}$ (7)
@@
$L_{t}=\min_{0\leq l<r\leq m}L_{t}(r,l)\ \ \text{and}\ \ \bar{L}\left(N\right)=\frac{\sum_{t=1}^{N}\omega_{t}\alpha_{t}^{2}L_{t}\mathbf{1}_{\mathcal{E}(t)}\left\|x^{t,r_{t}}-x^{t,l_{t}}\right\|^{2}/2}{\sum_{t=1}^{N}\mathbf{1}_{\mathcal{E}(t)}\left\|x^{t,r_{t}}-x^{t,l_{t}}\right\|^{2}}\ ,$ (8)
@@
Suppose function $f$ is convex, $\left(k,L\right)$-PWS and the sequence $\left\{\hat{x}^{t}\right\}$ generated by Algorithm 2 satisfies $f\left(\hat{x}^{t}\right)\geq\tilde{l}$ for all $t\geq 1$ and that every subproblem (7) of Algorithm 2 is feasible. Let $x^{t}:=x^{t,m}$, and assume the weight $\left\{\omega_{t}\right\}$ satisfy $\omega_{t}\left(1-\alpha_{t}/2\right)=\omega_{t-1}$ for $t\geq 1$. Then the following properties hold:
@@
    Convergence Bound: There exists a constant $\bar{L}(t)$ (defined in (8)) such that
    $\omega_{t}\left[f(\hat{x}^{t})-\tilde{l}\right]-\omega_{1}\left(1-\frac{\alpha_{1}}{2}\right)\left[f(\hat{x}^{0})-\tilde{l}\right]\leq\bar{L}(t)\|x^{t}-\bar{y}\|^{2}\ .$ (9)
@@
$\left\|x^{t,l_{t}}-\bar{y}\right\|^{2}+\left\|x^{t,l_{t}}-x^{t,r_{t}}\right\|^{2}\leq\left\|x^{t,r_{t}}-\bar{y}\right\|^{2}\ .$ (10)
Next, we give the descent analysis of the sequence based on (10). It follows from the non-decreasing property in Part1 and (10) that
@@
$\sum_{t=1}^{N}\left\|x^{t,l_{t}}-x^{t,r_{t}}\right\|^{2}\overset{(a)}{=}\left\|x^{1,0}-\bar{y}\right\|^{2}+\sum_{t=1}^{N}\left\|x^{t,l_{t}}-x^{t,r_{t}}\right\|^{2}\leq\left\|x^{N,m}-\bar{y}\right\|^{2}\ ,$ (11)
@@
$\begin{split}\omega_{t}\left(f\left(\hat{x}^{t}\right)-\tilde{l}\right)&\leq\omega_{t}\left(1-\frac{\alpha_{t}}{2}\right)\left(f\left(\hat{x}^{t-1}\right)-\tilde{l}\right)+\frac{\omega_{t}L_{t}\mathbf{1}_{\mathcal{E}(t)}\alpha_{t}^{2}}{2}\left\|x^{t,r_{t}}-x^{t,l_{t}}\right\|^{2}\\ &\overset{(a)}{=}\omega_{t-1}\left(f\left(\hat{x}^{t-1}\right)-\tilde{l}\right)+\frac{\omega_{t}L_{t}\mathbf{1}_{\mathcal{E}(t)}\alpha_{t}^{2}}{2}\left\|x^{t,r_{t}}-x^{t,l_{t}}\right\|^{2}\ ,\\ \omega_{1}\left(f\left(\hat{x}^{1}\right)-\tilde{l}\right)&\leq\omega_{1}\left(1-\frac{\alpha_{1}}{2}\right)\left(f\left(\hat{x}^{0}\right)-\tilde{l}\right)+\frac{\omega_{1}L_{1}\mathbf{1}_{\mathcal{E}(1)}\alpha_{1}^{2}}{2}\left\|x^{1,r_{1}}-x^{1,l_{1}}\right\|^{2}\ .\end{split}$ (12)
Summing (12) from $t=1$ to $N$ yields
$\omega_{N}\left(f\left(\hat{x}^{N}\right)-\tilde{l}\right)\leq\omega_{1}\left(1-\frac{\alpha_{1}}{2}\right)\left(f\left(\hat{x}^{0}\right)-\tilde{l}\right)+\bar{L}\left(N\right)\sum_{t=1}^{N}\mathbf{1}_{\mathcal{E}(t)}\left\|x^{t,r_{t}}-x^{t,l_{t}}\right\|^{2}\ .$ (13)
Combining $\sum_{t=1}^{N}\mathbf{1}_{\mathcal{E}(t)}\left\|x^{t,r_{t}}-x^{t,l_{t}}\right\|^{2}\leq\sum_{t=1}^{N}\left\|x^{t,r_{t}}-x^{t,l_{t}}\right\|^{2}$ and (13) gives
@@
where $(a)$ is by (11). Hence, we complete our proof. ∎∎
@@
First, Algorithm 2 is a double-loop method. By updating the weighting parameter $\alpha_{t}$ and the reference iterate $\hat{x}^{t}$ only once per outer iteration, the convergence analysis can compare arbitrary pairs of inner-loop iterates, rather than only successive ones, by applying the three-point lemma (see (10)). This flexibility enables the algorithm to exploit regions with smaller local Lipschitz constants, thereby better adapting to the PWS.
@@
$\text{Option I}:\ \ X(t,i):=\left\{\cap_{j\in\left[1,i\right]}\underline{X}(t,j)\right\}\cap\bar{X}\left(t,0\right)\ .$ (14)
@@
$\text{Option II}:\ \ X\left(t,i\right):=\left\{\cap_{j\in\left[1,i\right]}\underline{X}(t,j)\right\}\cap\bar{X}\left(t-1,0\right)\cap\left\{\cap_{j\in\left[1,m\right]}\underline{X}(t-1,j)\right\}$ (15)
@@
$\omega_{N}\left(f\left(\hat{x}^{N}\right)-\tilde{l}\right)-\omega_{1}\left(1-\frac{\alpha_{1}}{2}\right)\left(f\left(\hat{x}^{1}\right)-\tilde{l}\right)\leq\bar{L}\left(N\right)D_{X}^{2}\ ,$ (16)

### v2

@@ 1 Introduction
$f^{*}:=\min_{x\in X}f(x).$ (1.1)
Under the assumption that the objective $f$ satisfies the quadratic growth (QG) condition (see Definition 2), we investigate the oracle complexity of first-order methods for finding an $\epsilon$-optimal solution, defined as the theoretical number of (sub)-gradient evaluations required to reach a point $x$ such that $f(x)-f^{*}\leq\epsilon$. We focus on unknown piecewise smooth (PWS) functions: the domain $X$ admits a finite partition into regions on each of which the function is smooth, but the partition is not known a priori (Definition 1). Such functions arise, for example, in quadratic Model Predictive Control (MPC) [42] and two-stage stochastic linear programming (SLP) [16], where those regions are determined by the active constraints and the optimal dual solution in the second-stage problem, respectively, and hence are not available a priori. They also arise in ReLU networks [59], where regions are determined by the composition of activation functions. As network depth increases, these pieces become more difficult to characterize a priori.
@@
$\hat{L}_{k}=\frac{2\left[f(x^{k+1})-f(x^{k})-\left\langle f^{\prime}(x^{k}),x^{k+1}-x^{k}\right\rangle\right]}{\left\|x^{k+1}-x^{k}\right\|^{2}}.$ (1.2)
@@
Example 1.1: Average vs. worst-case empirical Lipschitz smoothness constant In the PWS setting, the gradient is continuous within each smooth region but may be discontinuous across their boundaries. This discontinuity can cause the empirical Lipschitz smoothness constant $\hat{L}_{k}$ (1.2) to spike when iterates $x^{k}$ and $x^{k+1}$ belong to different regions. Since the worst-case empirical Lipschitz smoothness constant $L_{\max}$ is determined by these spikes, it can be significantly larger than the average $L_{\text{avg}}$. To illustrate this, consider the iterates generated by applying the Polyak step size to the following one-dimensional objective function with two smooth regions, $\{x\geq 0\}$ and $\{x<0\}$: $\min_{x\in\left[-M,+\infty\right]}f(x;i)=(ix+1)\cdot\mathbf{1}_{\left\{x\geq 0\right\}}+e^{x}\cdot\mathbf{1}_{\left\{x<0\right\}}.$ The objective is parameterized by $i\in\mathbb{N}_{+}$ and $M\in\mathbb{R}$. Starting from an initial point $x^{0}=1/i$, we obtain $L_{\max}=i^{2}e^{-1/i}/2$, while $L_{\text{avg}}\rightarrow e^{-M}$ as the method progresses. Consequently, the ratio $L_{\max}/L_{\text{avg}}$ can be arbitrarily large for certain problem instances. Figure 2 illustrates how these two Lipschitz constants evolve over iterations.
@@
    Optimal convergence rate for PWS optimization. We introduce an accelerated bundle-level method termed Apex (Accelerated Prox-Level method for Exploring piecewise-smooth structure) in Section 3. We prove that with an adequate bundle of cutting planes, Apex achieves an optimal iteration complexity of $O\left(\kappa\sqrt{L/\mu^{*}}\log(1/\epsilon)\right)$ to compute an $\epsilon$-optimal solution for objectives satisfying $(\kappa,L)$-PWS (Definition 1) and $\mu^{*}$-QG (Definition 2) conditions. This result significantly improves upon the $O\left(\kappa^{2}L/\mu^{*}\log(1/\epsilon)\right)$ complexity reported in [74]. Our theoretical advancements translate to superior practical performance, as demonstrated in Figure 2. Furthermore, in Section 7, we establish the first lower complexity bound for PWS optimization, defining the theoretical benchmark for optimality. Consequently, Apex is the first algorithm to achieve an accelerated, provably optimal rate for piecewise-smooth optimization.
@@
    In Section 4, we introduce an accelerated bundle-level (BL) framework capable of constructing the W-certificate (see Definition 3) with significantly higher accuracy than previous methods [74]. While prior certification techniques could only bound the optimality gap within a factor of $O(L/\mu)$, our approach improves this bound to a constant factor, $O(1)$. This enhancement allows the certificate to serve as a robust proxy for the true optimality gap. Consequently, the certificate becomes a critical tool for dynamically updating the target level (an estimate of the optimal objective value, $f^{*}$), a essential parameter for the bundle-level method. Leveraging this accurate dynamic update, we design a new parameter-free algorithm, as detailed in our next contribution.
@@ 1.1 Related Work
$\frac{1}{2L}\left\|\nabla f\left(\bar{y}\right)\right\|^{2}\leq f\left(\bar{y}\right)-f^{*}\leq\frac{1}{2\mu}\left\|\nabla f\left(\bar{y}\right)\right\|^{2}\ .$ (1.3)
@@
For nonsmooth but Lipschitz functions, verifiability is preserved, but the strong guarantees in (1.3) no longer hold. By Rademacher’s theorem, Lipschitz functions are differentiable almost everywhere, which motivates defining the subdifferential as $\partial f(x):=\operatornamewithlimits{conv}\left\{g:g=\lim_{t\to\infty}f^{\prime}(x^{t}),x^{t}\to x\right\}$. With this definition, a point $x$ is stationary if $0\in\partial f(x)$. More generally, $x$ is a $(\delta,\varepsilon)$ near-approximate stationary point if $\left\|x-x^{\prime}\right\|\leq\varepsilon$ for some $\varepsilon$-stationary point $x^{\prime}$. While Tian and So [69] introduce the first oracle-polynomial-time algorithm for detecting near-approximate stationary points in piecewise affine functions, for general Lipschitz and bounded-below functions, finding such points can not be guaranteed in polynomial time [41]. As a relaxation, recent work [72, 41] studies $(\delta,\varepsilon)$-Goldstein stationarity, which considers the convex hull of subgradients in a $\delta$-neighborhood. Given a finite sample $\mathcal{P}=\left\{x^{i}\right\}_{i=1}^{m}\subseteq\mathcal{B}(x;\delta)$, one can test $(\delta,\varepsilon)$-Goldstein stationarity by: $\min_{\lambda\in\Delta_{+}}\left\|\sum_{i=1}^{m}\lambda_{i}f^{\prime}(x^{i})\right\|$, where $\Delta_{+}$ is the probability simplex. Approximate Goldstein stationarity is thus computationally tractable, though it does not imply bounds on the optimality gap. On the other hand, for structured problems such as weakly convex functions, the Moreau stationary certificate is widely used. The Moreau envelope [58] is defined as: $f_{\lambda}(x):=\min_{y}\left\{f(y)+\lambda\left\|y-x\right\|^{2}/2\right\}$, which is smooth when $\lambda$ exceeds the weak convexity modulus, with $\nabla f_{\lambda}(x)=\lambda\left(x-x_{\lambda}\right)$ and $x_{\lambda}=\operatornamewithlimits{argmin}_{y}\left\{f(y)+\lambda\left\|y-x\right\|^{2}/2\right\}$. The main challenge is that computing $x_{\lambda}$ exactly is as hard as solving the original problem [19]. Approximating Moreau stationary [6] is possible, but they require an approximation $\hat{x}_{\lambda}$ with $\left\|\hat{x}_{\lambda}-{x}_{\lambda}\right\|$ sufficiently small, which in turn demands an additional termination certificate for the proximal subproblem. Additionally, Nemirovski et al. [60] developed an accuracy certificate framework for convex optimization, variational inequalities with monotone operators, and convex Nash equilibrium problems, showing how such certificates can be constructed via ellipsoid or cutting-plane methods. Nevertheless, this framework has two crucial limitations: it only applies to bounded domains, and it does not accommodate error bounds such as the QG property. As a result, it produces only sublinear convergence rates for certificate generation.
@@ 2 Preliminaries
Definition 1 ($(k,L)$-Piecewise smoothness, $\left(k,L\right)$-PWS).
@@
Definition 2 ($\mu^{*}$-Quadratic growth, $\mu^{*}$-QG).
@@
$\begin{aligned} f(x^{t+1})\leq&\ell_{f}(x^{t+1};x^{t})+{\hat{L}_{t}}\left\|x^{t+1}-x^{t}\right\|^{2}/2\\ \leq&f^{*}+{\hat{L}_{t}}(\left\|x^{t}-x^{*}\right\|^{2}-\left\|x^{t+1}-x^{*}\right\|^{2})/2\end{aligned}\ ,$ (2.1)
@@
$\begin{aligned} f(x^{t+i})\leq&\ell_{f}(x^{t+i};x^{t})+{\hat{L}_{t,i}}\left\|x^{t+i}-x^{t}\right\|^{2}/2\\ \leq&f^{*}+i\cdot{\hat{L}_{t,i}}\left(\left\|x^{t}-x^{*}\right\|^{2}-\left\|x^{t+i}-x^{*}\right\|^{2}\right)/2\end{aligned}\ ,\forall i\in\left[m\right]\ ,$ (2.2)
@@
$\begin{split}\underline{x}^{t}&=\left(1-\alpha_{t}\right)\bar{x}^{t-1}+\alpha_{t}x^{t-1}\\ x^{t}&=\operatornamewithlimits{argmin}_{x\in X}\left\|x-x^{t-1}\right\|^{2}\ ,\ \operatornamewithlimits{s.t.}\ \ell_{f}\left(x;\underline{x}^{i}\right)\leq\tilde{l}\ ,\ i\in\left[t\right]\ ,\\ \bar{x}^{t}&=\left(1-\alpha_{t}\right)\bar{x}^{t-1}+\alpha_{t}x^{t}\end{split}$ (ABL)
@@
$f(\bar{x}^{t})\leq$ $\ell_{f}(\bar{x}^{t};\underline{x}^{t})+{\hat{L}_{t}}\left\|\bar{x}^{t}-\underline{x}^{t}\right\|^{2}/2$ (2.3)
$\leq$ $\alpha_{t}f^{*}+(1-\alpha_{t})f(\bar{x}^{t-1})+{\hat{L}_{t}}\alpha_{t}^{2}\left(\left\|x^{t}-\bar{y}\right\|^{2}-\left\|x^{t-1}-\bar{y}\right\|^{2}\right)/2$
where $\hat{L}_{t}$ is the empirical Lipschitz smoothness constant determined between $\bar{x}^{t}$ and $\underline{x}^{t}$. A key structural relationship, $\bar{x}^{t}-\underline{x}^{t}=\alpha_{t}(x^{t}-x^{t-1})$, forms the backbone of the telescoping arguments in the convergence proofs of ABL and APL. However, this identity is valid only for adjacent pairs of iterates and fails to generalize to non-adjacent sequences, such as those arising in the BL method discussed in (2.2). This limitation becomes especially pronounced when analyzing PWS functions, where telescoping across non-adjacent points is required.
@@ 3 Accelerated Prox-level method for piecewise smooth optimization
1:Initialize: $\text{{Flag}}=\text{{\text{False}}},x^{t+1,0}=x^{t,0}$
@@
$x^{t,i}\leftarrow\operatornamewithlimits{argmin}_{x\in X(t,i)}\left\|x-\bar{y}\right\|^{2}$ (3.1)
@@
$L_{t}=\min_{0\leq l<r\leq m}L_{t}(r,l)\ \ \text{and}\ \ \bar{L}\left(N\right)=\frac{\sum_{t=1}^{N}\omega_{t}\alpha_{t}^{2}L_{t}\mathbf{1}_{\mathcal{E}(t)}\left\|x^{t,r_{t}}-x^{t,l_{t}}\right\|^{2}/2}{\sum_{t=1}^{N}\mathbf{1}_{\mathcal{E}(t)}\left\|x^{t,r_{t}}-x^{t,l_{t}}\right\|^{2}}\ ,$ (3.2)
@@
Suppose function $f$ is convex, $\left(k,L\right)$-PWS and the sequence $\left\{\hat{x}^{t}\right\}$ generated by Algorithm 2 satisfies $f\left(\hat{x}^{t}\right)\geq\tilde{l}$ for all $t\geq 1$ and that every subproblem (3.1) of Algorithm 2 is feasible. Let $x^{t}:=x^{t,m}$, and assume the weight $\left\{\omega_{t}\right\}$ satisfy $\omega_{t}\left(1-\alpha_{t}/2\right)=\omega_{t-1}$ for $t\geq 1$. Then the following properties hold:
@@
    Convergence Bound: There exists a constant $\bar{L}(t)$ (defined in (3.2)) such that
    $\omega_{t}\left[f(\hat{x}^{t})-\tilde{l}\right]-\omega_{1}\left(1-\frac{\alpha_{1}}{2}\right)\left[f(\hat{x}^{0})-\tilde{l}\right]\leq\bar{L}(t)\|x^{t}-\bar{y}\|^{2}\ .$ (3.3)
@@
$\left\|x^{t,l_{t}}-\bar{y}\right\|^{2}+\left\|x^{t,l_{t}}-x^{t,r_{t}}\right\|^{2}\leq\left\|x^{t,r_{t}}-\bar{y}\right\|^{2}\ .$ (3.4)
Next, we give the descent analysis of the sequence based on (3.4). It follows from the non-decreasing property in Part1 and (3.4) that
@@
$\sum_{t=1}^{N}\left\|x^{t,l_{t}}-x^{t,r_{t}}\right\|^{2}\overset{(a)}{=}\left\|x^{1,0}-\bar{y}\right\|^{2}+\sum_{t=1}^{N}\left\|x^{t,l_{t}}-x^{t,r_{t}}\right\|^{2}\leq\left\|x^{N,m}-\bar{y}\right\|^{2}\ ,$ (3.5)
@@
$\begin{split}\omega_{t}\left(f\left(\hat{x}^{t}\right)-\tilde{l}\right)&\leq\omega_{t}\left(1-\frac{\alpha_{t}}{2}\right)\left(f\left(\hat{x}^{t-1}\right)-\tilde{l}\right)+\frac{\omega_{t}L_{t}\mathbf{1}_{\mathcal{E}(t)}\alpha_{t}^{2}}{2}\left\|x^{t,r_{t}}-x^{t,l_{t}}\right\|^{2}\\ &\overset{(a)}{=}\omega_{t-1}\left(f\left(\hat{x}^{t-1}\right)-\tilde{l}\right)+\frac{\omega_{t}L_{t}\mathbf{1}_{\mathcal{E}(t)}\alpha_{t}^{2}}{2}\left\|x^{t,r_{t}}-x^{t,l_{t}}\right\|^{2}\ ,\\ \omega_{1}\left(f\left(\hat{x}^{1}\right)-\tilde{l}\right)&\leq\omega_{1}\left(1-\frac{\alpha_{1}}{2}\right)\left(f\left(\hat{x}^{0}\right)-\tilde{l}\right)+\frac{\omega_{1}L_{1}\mathbf{1}_{\mathcal{E}(1)}\alpha_{1}^{2}}{2}\left\|x^{1,r_{1}}-x^{1,l_{1}}\right\|^{2}\ ,\end{split}$ (3.6)
where $(a)$ holds by the definition of $\omega_{t}$. Summing (3.6) from $t=1$ to $N$ yields
$\omega_{N}\left(f\left(\hat{x}^{N}\right)-\tilde{l}\right)\leq\omega_{1}\left(1-\frac{\alpha_{1}}{2}\right)\left(f\left(\hat{x}^{0}\right)-\tilde{l}\right)+\bar{L}\left(N\right)\sum_{t=1}^{N}\mathbf{1}_{\mathcal{E}(t)}\left\|x^{t,r_{t}}-x^{t,l_{t}}\right\|^{2}\ .$ (3.7)
Combining $\sum_{t=1}^{N}\mathbf{1}_{\mathcal{E}(t)}\left\|x^{t,r_{t}}-x^{t,l_{t}}\right\|^{2}\leq\sum_{t=1}^{N}\left\|x^{t,r_{t}}-x^{t,l_{t}}\right\|^{2}$ and (3.7) gives
@@
where $(a)$ is by (3.5). Hence, we complete our proof.
@@
First, Algorithm 2 is a double-loop method. By updating the weighting parameter $\alpha_{t}$ and the reference iterate $\hat{x}^{t}$ only once per outer iteration, the convergence analysis can compare arbitrary pairs of inner-loop iterates, rather than only successive ones, by applying the three-point lemma (see (3.4)). This flexibility enables the algorithm to exploit regions with smaller local Lipschitz constants, thereby better adapting to the PWS.
@@
$\text{Option I}:\ \ X(t,i):=\left\{\cap_{j\in\left[1,i\right]}\underline{X}(t,j)\right\}\cap\bar{X}\left(t,0\right)\ .$ (3.8)
@@
$\text{Option II}:\ \ X\left(t,i\right):=\left\{\cap_{j\in\left[1,i\right]}\underline{X}(t,j)\right\}\cap\bar{X}\left(t-1,0\right)\cap\left\{\cap_{j\in\left[1,m\right]}\underline{X}(t-1,j)\right\}$ (3.9)
@@
$\omega_{N}\left(f\left(\hat{x}^{N}\right)-\tilde{l}\right)-\omega_{1}\left(1-\frac{\alpha_{1}}{2}\right)\left(f\left(\hat{x}^{1}\right)-\tilde{l}\right)\leq\bar{L}\left(N\right)D_{X}^{2}\ ,$ (3.10)